OpenStudy (anonymous):

Look i know this is not what this site is for but i really need some help i am way behind on my credits and am taking online classes i have taken this one test multiple times and can't seem to pass it i am down to my last attempt to pass it and i am getting really frustrated. i would appreciate any help or answers u can give me.

OpenStudy (anonymous):

A college student is looking at her monthly checking account records. On September 1, 2008, her checking account held a balance of $1,050. At the end of March 2009, her checking account held a balance of $800. What is the monthly rate of change for the student’s checking account over the period September 1, 2008 through March 31, 2009? A. -1850/7 B. -250/7 C. 250/7 D. 1850/7

OpenStudy (anonymous):

A line passes through the point [0,-3/2] and has a slope of 1/2. What is the equation of the line? A. y= 1/2x-3/2 B. y=1/2x+3/2 C. y=-1/2x-3/2 D. y=1/2x-1/2

OpenStudy (anonymous):

Which equation is equivalent to 3x– 4y= 8? A. y=3/4x+2 B. y=-3/4x+2 C. y=-3/4x-2 D. y=3/4x-2

OpenStudy (anonymous):

Which equation is equivalent to 2x+ 6y= 12? A. y=1/3x+2 B. y=1/3x-2 C. y=-1/3x+2 D. y=-1/3x-2

OpenStudy (anonymous):

A line passes through the point (0, –1) and has a slope of -2/3. What is the equation of the line? A. y=-2/3x-1 B. y=3/2x+1 C. y=2/3x+1 D. y=-2/3x+1

OpenStudy (anonymous):

Which equation is equivalent to 3x– 4y= 8? Which equation is equivalent to 2x+ 6y= 12? Solve each equation above for y. Move the x term to the RHS and then divide both sides by the coefficient of x.